# RSTeller: Scaling Up Visual Language Modeling in Remote Sensing with Rich Linguistic Semantics from Openly Available Data and Large Language Models

### v1

@@ 5 Discussions
First, the results in domain data scaling (Section 4.3) confirm that increasing the volume of domain-specific data generally enhances the model performance across various benchmarks. This is evidenced by the consistent improvement in both zero-shot classification and image retrieval tasks as more domain data is included. However, the degree of improvement varies among different benchmarks, indicating that the effectiveness of domain-specific data augmentation is influenced by the nature of the benchmark dataset. For instance, benchmarks like PatternNet and AID showed significant performance gains, whereas others like EuroSAT exhibited less pronounced improvements or even anomalies in certain cases.
One notable observation is that the initial performance degradation observed in some benchmarks, such as the AID and RSI-CB, when domain-specific data is first introduced. Despite this initial setback, continuous scaling of the domain data eventually leads to performance improvements, suggesting that the models require a certain threshold of domain-specific data to effectively learn and generalize. This indicates the importance of not being discouraged by the initial performance drops and continuing to expand the dataset for long-term gains.
@@
Finally, the experiments in testing the RSTeller dataset (Section 4.3) revealed that the starting checkpoint plays a critical role in the effectiveness of continual pre-training. Models pre-trained on DataComp outperformed those pre-trained on other datasets in most cases, suggesting that selecting a robust initial checkpoint can lead to better results in continual pre-training. This finding emphasizes the importance of choosing the right pre-trained model as a starting point for further training with the RSTeller dataset.
@@ 6 Limitations of the Proposed Work
First, the captioning tasks within our dataset generation workflow are simplistic, typically focusing on a single element, which might not capture all pertinent details in the images. For instance, image patch 4 in Fig. 7 is captioned solely by its depiction of an industrial complex, omitting other significant elements like cars and the parking lot within. Similarly, image patch 10 focuses only on a single solar panel within a larger solar power plant, potentially misleading in terms of scale, location, and quantity. Future enhancements could include more intricate captioning tasks to encompass a broader range of information within the images.
@@ 7 Conclusion
In this paper, we have meticulously developed and presented an automated dataset generation workflow that enables researchers to efficiently gather and utilize multimodal data for RS applications on a large scale. This workflow optimizes the use of existing, openly available remote-sensing data and employs an LLM to transform raw data into a high-quality, usable dataset for training VLMs in the remote sensing domain. Central to our workflow is the introduction of RSTeller, a multimodal dataset comprising over 2.5 million high-quality image-text pairs with diverse and informative captions, suitable for a wide range of research tasks. Utilizing this dataset for training RS VLMs not only facilitates access to high-quality data for researchers but also reduces the environmental impact associated with generating and maintaining such a dataset. Moreover, we highlight the potential for significant advancements in the field of RS VLMs through the continual enrichment of RS-related multimodal data. We envision future research expanding these datasets to a larger scale, thus enabling the training of more robust VLMs and enhancing the potential prosperity achievable in this field.
@@
During the preparation of this work the authors used ChatGPT in order to improve readability and language. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication
@@ Acknowledgments
This study was supported by the National Natural Science Foundation of China (Nos. 62101416, 62301405), the Natural Science Basic Research Program of Shaanxi, China (2022JQ-708), and Fundamental Research Funds for the Central Universities, China.

### v4

@@ 4.3 Dataset Effectiveness
4.4 Ablation Study on Training Data
To assess the necessity and effectiveness of the individual components in generating the RSTeller dataset, we conducted an ablation study. This study revisits the previous experiment with different settings, where we systematically removed one component at a time. The baseline for this experiment is based on the ViT-L/14 model, pre-trained on WIT, followed by continual pre-training on both RSTeller and LAION10M as discussed in Section 4.3. The results of this ablation study are shown in Table 11, with the first row displaying the baseline results and the subsequent rows showing the performance differences for each setting. Positive differences are highlighted in blue.
In the first experiment, we examined the effect of excluding the LAION10M dataset during the continual pre-training process. Removing this common knowledge resulted in a performance drop across most benchmarks, not only on ImageNet-1K but also on several remote sensing tasks. This suggests that certain common knowledge, present in LAION10M, contributes to the model’s ability to generalize well to new domains. Interestingly, the EuroSAT classification task showed a slight improvement. Upon reviewing the confusion matrices, we found that the performance gain primarily came from improved classification of the “permanent crop” and “residential” classes. This improvement suggests that there might be some noise within the LAION10M dataset that could negatively impact the model’s performance on specific classes in the EuroSAT benchmark.
The RSTeller dataset includes named entities such as “name” or “tiger:county,” which may not directly correspond to visual features in remote sensing images and could potentially introduce noise. To assess their impact, we removed named entities using the spaCy library (Honnibal et al., 2020), and retrained the model on this modified dataset. The results revealed a negligible drop in performance, indicating that these entities are not detrimental to model performance. While some named entities, such as “RE Kent South” from Fig. 3(a), may appear irrelevant to visual content, others, like “Stonycreek River” from Fig. 3(d), provide descriptive terms (e.g., “river”) closely tied to visual features, contributing valuable contextual information during pretraining. Furthermore, named entities may enable broader applications, such as retrieval tasks where specific names or geo-elements are queried. Given their negligible negative impact and potential utility, we conclude that retaining named entities in the dataset is beneficial.
A key contribution of our work is leveraging LLMs to interpret OSM data related to remote sensing images and generate corresponding captions. However, LLMs may introduce hallucinations when inferring the meaning of OSM tags. To investigate the importance of LLM interpretation, we conducted the third experiment by excluding LLMs from the caption generation process. In this experiment, captions were generated using a predefined template, namely “A remote sensing image of [tag1]; [tag2]; …,” formatted with the same filtered OSM tags as in the RSTeller dataset. The results revealed a significant drop in overall performance, underscoring the necessity of including LLMs in the caption generation process. However, we observed that certain benchmarks, including fMoW in classification and RSICD in retrieval, showed improvements when LLM interpretation was excluded. These datasets contain finer-grained classes or more specific content, which may have been obscured by the LLM interpretation. This finding suggests that further research is needed to refine the process for generating more precise captions.
Finally, we evaluated the impact of Task 3, which involves using LLMs to augment captions by revising them in different styles. In this experiment, we removed the Task 3 captions from the dataset. The results demonstrated a notable decline in performance, highlighting the value of this component in the caption generation process. While the performance gain from Task 3 was not large, the task is relatively simple and resource-efficient, requiring fewer tokens for inference. Additionally, revising captions into various styles makes the data synthesis process more versatile and accessible to researchers, especially when raw remote sensing images or OSM data are not readily available. Thus, Task 3 shows promise as a useful tool for enhancing existing multimodal datasets in the remote sensing domain.
4.5 Model Performance wrt. Domain Data Scale
Fig. 11: Relationship between domain data scale and zero-shot classification performance. Left: A ViT-B/32 model. Right: A ViT-L/14 CLIP model. Both models are initialized with weight trained on WIT. Scaling domain data consistently improves the performance of the model on the validation dataset.
Fig. 12: Relationship between domain data scale and zero-shot image retrieval performance. Left: A ViT-B/32 model. Right: A ViT-L/14 CLIP model. Both models are initialized with weight trained on WIT. Scaling domain data consistently improves the performance of the model on the validation dataset.
This experiment was designed to investigate whether scaling the domain-specific data enhances model performance. To this end, we amalgamated our RSTeller dataset with the SkyScript dataset to create a robust RS domain dataset comprising approximately 5.2 million image-text pairs. This extensive dataset allows for a detailed analysis of the model’s performance as it relates to variations in domain data volume. We partitioned the data into four equal subsets, each representing about one-fourth of the total data, and ensured no overlap between the subsets. In our experiments, we incrementally incorporated these subsets into the training process, thereby systematically increasing the domain data scale. The ViT-B/32 and ViT-L/14 models pre-trained on WIT were used as the starting point for the continual pre-training. The models were then trained on approximately 25%, 50%, 75%, and 100% of the domain data, respectively.
Most benchmarks show improvement when domain data is scaled up, as indicated by the negative slopes of the error rates for both models ($k_{Average}=-0.50$ for ViT-B/32 and $k_{Average}=-0.63$ for ViT-L/14). Datasets such as PatternNet, RESISC, and EuroSAT exhibit substantial improvements, with linear fit slopes of $k_{PatternNet}=-1.14$, $k_{RESISC}=-0.64$ for the ViT-B/32 model, and $k_{EuroSAT}=-1.50$ for the ViT-L/14 model. These results highlight the positive impact of enlarging the domain-specific dataset on the model’s classification accuracy.
However, an unexpected trend is observed with the RSI-CB dataset for the ViT-L/14 model, where top-1 classification error increases as the domain data scale grows ($k_{RSI-CB}=0.90$). Analyzing the model’s predictions revealed that certain classes were more frequently misclassified as more RS domain data were incorporated. For instance, “bare land” was mistaken for “desert,” “green farmland” for “stream,” and “mangrove” or “sparse forest” for “forest,” while “resident” was misclassified as “avenue.” These classes often share similar visual features or appear in close proximity within a single image, suggesting potential bias due to granularity and the choice of OSM elements for captioning.
For some benchmarks, such as AID and RESISC evaluated on ViT-L/14, there is an initial performance degradation when domain-specific data is first introduced. In extreme cases, even with all the domain data used for training, the fine-tuned model performance still does not surpass the baseline (e.g., AID for ViT-L/14). Nevertheless, the continual inclusion of domain data leads to consistent performance improvement. For instance, the RESISC benchmark shows a steady decline in error rates despite the initial degradation, as indicated by the slopes ($k_{RESISC}=-0.89$ for ViT-L/14). This trend suggests that further scaling of domain data could potentially surpass the baseline performance of models pre-trained on generic web-based data.
The varied slopes of the linear fits also highlight the differential impact of domain data across different benchmarks. For example, in the ViT-B/32 experiment, the significant negative slope observed for PatternNet ($k_{PatternNet}=-1.14$) compared to the minimal impact on Million-AID ($k_{Million-AID}=-0.00$) suggests that the effectiveness of domain-specific data augmentation varies across different datasets. This variation indicates inherent differences in the benchmarks themselves and possibly the domain specificity of the data they contain.
The experimental results for zero-shot image retrieval tasks, illustrated in Fig. 12, reveal trends consistent with those observed in the classification experiments. The x-axis represents the volume of RS-related domain data used for continual pre-training, while the y-axis indicates the evaluation metric, which is the average recall subtracted from 100.
For both the ViT-B/32 and ViT-L/14 models, there is a general improvement in performance with increasing domain data, as indicated by the negative slopes of the linear fits ($k_{Average}=-0.17$ for ViT-B/32 and $k_{Average}=-0.84$ for ViT-L/14). This trend is evident across multiple benchmarks, highlighting the benefits of enlarging domain-specific datasets for image retrieval tasks.
The experimental results confirm that increasing domain-specific data generally enhances model performance in the field of remote sensing. Despite the initial degradation observed in specific benchmarks, continuous improvement with increased data volume underscores the potential for further advancements. These findings validate the significance of this work, highlighting the proposed automated workflow’s efficiency in generating larger datasets at a low cost. This capability offers substantial benefits to the research community, particularly for training VLMs related to remote sensing.
@@ 5 Discussions
The experiments in testing the RSTeller dataset (Section 4.3) revealed that the starting checkpoint plays a critical role in the effectiveness of continual pre-training. Models pre-trained on DataComp outperformed those pre-trained on other datasets in most cases, suggesting that selecting a robust initial checkpoint can lead to better results in continual pre-training.
The results in domain data scaling (Section 4.5) confirm that increasing the volume of domain-specific data generally enhances the model performance across various benchmarks. This is evidenced by the consistent improvement in both zero-shot classification and image retrieval tasks as more domain data are included. However, the degree of improvement varies among different benchmarks, indicating that the effectiveness of domain-specific data augmentation is influenced by the nature of the benchmark dataset. For instance, benchmarks like PatternNet and EuroSAT showed significant performance gains, whereas others like RSI-CB exhibited less pronounced improvements or even anomalies in certain cases.
One notable observation is that the initial performance degradation observed in some benchmarks, such as the AID and RESISC, when domain-specific data are first introduced. Despite this initial setback, continuous scaling of the domain data eventually leads to performance improvements, suggesting that the models require a certain threshold of domain-specific data to effectively learn and generalize. This indicates the importance of not being discouraged by the initial performance drops and continuing to expand the dataset for long-term gains.
@@ 6 Limitations of the Proposed Work
First, the captioning tasks within our dataset generation workflow are simplistic, typically focusing on a single element, which might not capture all pertinent details in the images. For instance, Fig. 3(b) is captioned solely by its depiction of farmlands, omitting other significant elements like houses and the road nearby.
Additionally, while our efforts—such as task simplification, tag filtering, wiki-based tag interpretation, and caption refinement—substantially mitigate hallucinations, some residual errors may persist due to the inherent limitations of LLMs. These limitations can occasionally result in speculative or inaccurate outputs. Future work could explore advanced strategies to further suppress hallucinations, including leveraging vision-language models to rate and filter generated captions or integrating uncertainty quantification techniques to flag potentially unreliable outputs.
@@ 7 Conclusion
In this paper, we have meticulously developed and presented an automated dataset generation workflow that enables researchers to efficiently gather and utilize multimodal data for RS applications on a large scale. This workflow optimizes the use of existing, openly available remote-sensing data and employs an LLM to transform raw data into a high-quality, usable dataset for training VLMs in the remote sensing domain. Central to our workflow is the introduction of RSTeller, a multimodal dataset comprising over 2.6 million high-quality image-text pairs with diverse and informative captions, suitable for a wide range of research tasks. Utilizing this dataset for training RS VLMs not only facilitates access to high-quality data for researchers but also reduces the environmental impact associated with generating and maintaining such a dataset. Moreover, we highlight the potential for significant advancements in the field of RS VLMs through the continual enrichment of RS-related multimodal data. We envision future research expanding these datasets to a larger scale, thus enabling the training of more robust VLMs and enhancing the potential prosperity achievable in this field.
@@
During the preparation of this work the authors used ChatGPT in order to improve readability and language. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Acknowledgments
This study was supported by the National Natural Science Foundation of China (Grant Nos. 62476205 and 62301405), and the Fundamental Research Funds for the Central Universities (ZYTS25152).
Appendix A Regular Expression for Tag Filtering
This section presents the regular expression patterns used for tag filtering in the RSTeller dataset.
patterns = [
"^(tiger:(?!seperated|county).*|massgis:.*|nysgissam.*)",
"^gnis:(f.*|created|.*id|edited)",
"^(addr:flats|ref.*|FMMP_.*|created_by|source)",
".*(wikidata|wikipedia|address|postcode|housenumber|phone|website).*",
"^[Nn][Hh][Dd](?![:][Ff][Tt][Yy][Pp][Ee]).*"